CH4+2H2O+heat <--> CO2+4H2

A test reaction is done starting with only methane and excess water in the autoclave. If the initial concentration of methane is 0.112 mol/L, and the methane concentration at equilibrium is 0.017 mol / L, then the equilibrium concentration of hydrogen is?
(two digit answer)
1
Expert's answer
2020-06-02T14:09:05-0400

According to the reaction equation:

CH4+2H2O+heat "\rightarrow" CO2+4H2,

1 mol of methane produces 4 moles of hydrogen. The concentration of methane decreased by 0.112 - 0.017 = 0.095 M. As the initial concentration of hydrogen in the autoclave is zero, the concentration of hydrogen at the equilibrium is :

"[H_2] = 0+ 4(c_{CH_4}^0 - [CH_4]) = 4\u00b70.095=0.38" M.

Answer: the equilibrium concentration of hydrogen is 0.38 M.